a piper cub airplane has an avarage speed that is 30 mph faster than a cessna 150 airplane. it takes the cessna 5 hr to fly from fort lauderdale to atlanta, which it takes the piper cub only 4 hr to make the same trip. what is the avarage speed of each plane?
Answer by richwmiller(17219)

p=c+30
rt=d
5c=(c+30)*4
5c=4c+120
c=120
p=150
check
5*120=4*150
600=600
ok